OpenStudy (anonymous):

Can some please help me with this: Suppose f(3)=4, f(8)=8, f'(3)=2, f'(8)=9, and f'' is continuous. Find the value of the definite integral

OpenStudy (anonymous):

\[\int\limits_{3}^{8}\] xf''(x)

OpenStudy (anonymous):

I need this broken down in layman terms

OpenStudy (amistre64):

well, if we try to use this according to the definition of integration .. what can we get?

OpenStudy (amistre64):

\[f(x)=\int f'(x)dx\]is the basic understanding of an integral

OpenStudy (anonymous):

I don't understand what to do with the function values

OpenStudy (amistre64):

it would appear to me that the 3 to 8 is the limits of integration

OpenStudy (amistre64):

\[\int_{3}^{8}x\ f''(x) dx\] is what we have to work with and all we have to determine is what to do with the "x" part

OpenStudy (amistre64):

i might have to logout and back in to do this ...

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the math still aint processing from the latex on my end; you see normal equations on your end?

OpenStudy (anonymous):

yes I can see them.

OpenStudy (amistre64):

good :) well then, lets see if we can move ahead with this then i think we can apply integration by parts to this .. just a gut felling

OpenStudy (amistre64):

\[\int u dv =uv-\int v du\]

OpenStudy (anonymous):

ok so would i just use xf''(x) as my function to integrate?

OpenStudy (amistre64):

to wit: u = x v = f'(x) du = dx dv = f''(x) dx yes

OpenStudy (anonymous):

so when I integrate that is where the values that were given will be used?

OpenStudy (amistre64):

\[\int xf''(x)=xf'(x)-\int f'(x)dx\] \[\int xf''(x)=xf'(x)-f(x)\]

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

im sure that looks alot more inspiring on your end lol

OpenStudy (anonymous):

I shows up normal on my end

OpenStudy (amistre64):

(8 f'(8)) - f(8) ) - (3 f'(3)-f(3)) should be our results

OpenStudy (anonymous):

ok I didn't get that far but how did you get rid of the x?

OpenStudy (amistre64):

I didnt; i just used it in the integration by parts formula. there is no reason to get rid of it since the interval speaks in x to begin with

OpenStudy (anonymous):

ok i see know you used the limit that you are evaluating the integral?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

I got 53, is that correct?

OpenStudy (anonymous):

I see my error, I am about to recalculate it

OpenStudy (amistre64):

8*9 - 8 - (3*2-4) 72-8 - (6-4) 64 - 2 = 62

OpenStudy (amistre64):

i hope that makes more sense, I have to be heading out now. good luck :)

OpenStudy (anonymous):

I accidently used 8*8-9-)-(3*2-4). Thank you that was really helpful.